OpenStudy (aravindg):

a particle moving with uniform acceleration from A to B along a straight line has velocities v1 and v2 at A and B respectively.if C is midpoint between A and B determine velocity of particle at c

OpenStudy (nikvist):

\[v_2^2=v_1^2+2as\quad,\quad as=\frac{v_2^2-v_1^2}{2}\]\[v_C^2=v_1^2+as=v_1^2+\frac{v_2^2-v_1^2}{2}=\frac{v_2^2+v_1^2}{2}\]\[v_C=\sqrt{\frac{v_2^2+v_1^2}{2}}\]